OpenStudy (zzr0ck3r):

If \(X\) is a Banach space, then every absolutely convergent series is convergent. \(\mathit{pf:}\) Suppose s_n is absolutely convergent. For any \(m

OpenStudy (zzr0ck3r):

@eliassaab :)

OpenStudy (zzr0ck3r):

is it true that every convergent sequence is Cauchy in a complete normed space?

OpenStudy (zzr0ck3r):

if that is true then we choose n,m>N (the N for absolute convergence) and we are done...I think.....

OpenStudy (zzr0ck3r):

Ok I think that is it, but ill leave it up for you to confirm that is the point of the proof:)

OpenStudy (anonymous):

Since \[ \sum_{k=0}^\infty||x_k|| \] is convergent, for each \( \epsilon>0 \)then there is N, such that for n greater that N \[ \sum_{k=n}^\infty||x_k|| \le \epsilon \] That is what you need to finish your proof.

OpenStudy (zzr0ck3r):

and this is true because its Cauchy?

OpenStudy (anonymous):

Let \( \epsilon >0\) and let N> found above then \[ m>n> N \implies || s_m-s_n|| =||\sum_{k=n+1}^m x_k|| \le \sum_{k=n+1}^m||x_k|| \le \sum_{k=n+1}^\infty ||x_k|\le \epsilon \]

OpenStudy (anonymous):

Now s_n is Cauchy, so it is convergent in a Banach space, since it is complete.

OpenStudy (zzr0ck3r):

right, great. tnx again. Hibert spaces are next:)

OpenStudy (anonymous):

YW

OpenStudy (zzr0ck3r):

is the converse hard to prove?

OpenStudy (zzr0ck3r):

nm I found it.